# ValueScope: Unveiling Implicit Norms and Values via Return Potential Model of Social Interactions

### v1

@@ 1 Introduction
Social norms—the perceived, informal, and mostly unwritten rules that govern acceptable behaviors within a community—are foundational to understanding the dynamics of social interactions and shaping the community’s identity UNICEF (2021). Social values, in turn, are the deeper ideals and principles that a community aspires to uphold, guiding the creation and enforcement of these norms (McClintock, 1978). Social norms and values emerge organically through the interplay of behaviors Bicchieri et al. (2023) and are difficult to grasp without gaining experience of the community firsthand. This complexity poses challenges for new users to assimilate Lampe et al. (2014) and makes it difficult for automatic community moderation systems, which lack the nuanced understanding of in-community human moderators (Park et al., 2021).
@@
    Our work offers new insights into the dynamics of social norms, especially how they evolve over time. These findings have important scientific and practical implications for social scientists and community moderators, helping them identify norms that are likely to change and enabling proactive intervention.
@@ 2 Related Works
Norms in Online Communities   Computational studies have examined linguistic norms and semantic changes in online communities Lucy and Bamman (2021); Del Tredici and Fernández (2018); Kershaw et al. (2016); Danescu-Niculescu-Mizil et al. (2013); Hemphill and Otterbacher (2012); Del Tredici and Fernández (2017); Snoswell et al. (2023); Chancellor et al. (2018). However, these often focus narrowly on language use and neologisms, neglecting the wider spectrum of community values influenced by feedback. While some research has addressed explicit governance Chandrasekharan et al. (2018); Fiesler et al. (2018); Park et al. (2021) or qualitatively studied implicit norms Kasunic and Kaufman (2018); Shen and Rosé (2022), our approach fills the gap by (1) focusing on a range of implicit norms (e.g., formality and sarcasm) automatically selected through a generalizable norm induction process, and (2) analyzing collective community preference over behaviors along the selected norm dimensions to capture a comprehensive spectrum of community values, which can provide more fine-grained and objective measurement for alignment Bergman et al. (2024); Findeis et al. (2024).
@@ 3.1 The ValueScope Framework
Theoretical Background   Community members acquire social adeptness by learning unwritten rules, or implicit norms, that govern appropriate actions under various conditions, with feedback from others to guide their behaviors (Coutu, 1951; Zhang et al., 2023). The Return Potential Model (Jackson, 1966, RPM) quantifies these norms by mapping the return potential—expected (dis)approval—across different behaviors. Individuals in a community adjust their actions based on the learned mental model of return potential. We propose ValueScope, a computational framework that adapts RPM to analyze the expected community preference to behaviors with varying normness scales (i.e., conforming to a norm dimension to different extents), offering scalable insights into community values.
Problem Definition   Let $\mathcal{C}$ be communities, $\mathcal{A}$ be comments, and $\mathcal{D}$ be norm dimensions (e.g., sarcasm). For an arbitrary community $c\in\mathcal{C}$ and norm dimension $d\in\mathcal{D}$, ValueScope measures the normness scale $\Phi$ via the Normness Scale Predictor, $\Phi_{d}:\mathcal{A}\rightarrow\mathbb{R}$, and the community preference $\Psi$ via the Community Preference Predictor, $\Psi_{c}:\mathcal{A}\rightarrow\mathbb{R}$, of all $N$ comments in $c$: $\mathcal{A}_{c}$. [^2] Here, we define $\Phi_{d}$ and $\Psi_{c}$ to be the vanilla normness scale and community preference; to obtain the distilled scores derived in §3.2 and §3.3, we simply take the delta between two comments $(a_{i},a_{i}^{\prime})$ to get $\nabla\Phi_{d}:(\mathcal{A}\times\mathcal{A})\rightarrow\mathbb{R}=\Phi_{d}(a_{i}^{\prime})-\Phi_{d}(a_{i})$ and $\nabla\Psi_{c}:(\mathcal{A}\times\mathcal{A})\rightarrow\mathbb{R}=\Psi_{c}(a_{i}^{\prime})-\Psi_{c}(a_{i})$. For an arbitrary range of normness scales $\Phi_{d}^{i}:=[\phi_{d}^{\prime},\phi_{d}^{\prime\prime})$ (e.g., “somewhat sarcastic”), we take the set of comments $\mathcal{A}_{c,d}^{i}:=\{a_{i}|\Phi_{d}(a_{i})\in\Phi_{d}^{i}\}$ with normness scales in the given range, and let $N_{c,d}^{i}:=||\mathcal{A}_{c,d}^{i}||$ be the number of comments in this subset. We compute the community preference of these comments:
@@
Interpreting ValueScope   Via ValueScope, we quantitatively observe a number of features of the RPM model proposed in social science literature Jackson (1966); Nolan (2015); Linnan et al. (2005), enhancing our understanding of community values. Specifically, we use the point of maximum return—the highest point on the RPM curve—to locate the ideal normative behavior one should follow to maximize community preference, and the potential return difference—total positive feedback minus total negative feedback—to discover norm regulation strategies; i.e., whether the community tends to use reward or punishment to guide the formation and adaptation of its values.
@@ 3.2 Normness Scale Predictor (NSP)
Normness Distillation   The normness distillation stage addresses two key challenges. First, unlike survey-based social science studies, our approach observes normative behaviors post-hoc, lacking the opportunity to explore “alternative behaviors.” We attempt to recreate the “hypothetical conditions” proposed in Jackson (1966), in which the individual considers alternative options to maximize return (Zhang et al., 2023). We achieve this with a community language simulation module, which generate comments identical to the original, except for controlled variations in one norm dimension. We then apply the normness measurement module to quantify the normness scales of the transformed comments. E.g., for an original comment, “ty!,” we generate “thank you” by varying formality, and obtain formality scales of $0.2$ and $0.4$, respectively.
@@ 4.2.1 Normness Measurement
Data Preprocessing   Each topical group and norm dimension except for the verbosity dimension  [^4] Instead of training a verbosity scale classifier, we use the number of characters to intuitively measure verbosity. Using the character counts, we compute winrates in the range [0-1] to match the scale of other dimensions. has a dedicated classifier model, enabling comparisons across similar subreddits. Normness measurement relies on synthetic labels generated through stratified sampling and automatic labeling. During the sampling stage, comments are rated on a 5-point Likert scale by GPT-3.5 (Brown et al., 2020) to gauge normness (see Appendix D for the Likert scale details; Appendix F.1 for GPT-3.5 rating evaluation details). Then, 10 comments are sampled per scale point per subreddit, resulting in 150 comments per topic (200 for finance with 4 subreddits included). From these, 1,250 comment pairs are randomly selected to create binary synthetic labels using GPT-4  [^5] Cost estimations of the stratified sampling (GPT-3.5) and pairwise labeling (GPT-4) processes are in Appendix M. (OpenAI et al., 2024); we detail the GPT-4 prompt tuning and synthetic label evaluations in Appendix F.2. We train DeBERTa-base He et al. (2020) with the synthetic labels for each of the 4 topic groups and 5 norm dimensions with training details in Appendix G.1.
Gender Politics Science Finance
Acc. F1 Acc. F1 Acc. F1 Acc. F1
GPT-4 81.8 82.4 73.7 75.2 80.8 82.1 81.3 82.2
Val. 84.4 84.3 83.2 83.1 88.6 88.5 83.9 83.8
Test 78.1 77.9 74.3 74.2 83.0 83.0 78.1 78.1
Table 1: Normness Scale Predictor evaluation for each topic, averaged across norm dimensions: accuracy (Acc.) and F1 scores (F1). We evaluate GPT-4 generated labels against human annotations (GPT-4) and the binary classifier predictions against a held-out set with synthetic labels (Val.) and human annotations (Test).
Evaluation   To evaluate the quality of GPT-4 generated training labels and the NSP models, we curate a high-quality human annotation set of 450 samples for each norm dimension, where each sample is annotated by 3 annotators with an average inter-annotator agreement, measured by Fleiss’s kappa, of 0.58 (see Appendix E.1 for annotation details). Quantitative evaluation results of the GPT-4 generated synthetic labels and trained classifiers using human annotations and a held-out validation split of the synthetic labels are reported in Table 1, validating the quality of the NSP models. We report detailed results for GPT-4 labels in Appendix F.3 and for the NSP models in Appendix G.3.
@@ 4.2.2 Community Language Simulation
The norm distillation stage of NSP employs a community language simulation module to synthesize comments and control for norm variations.
Data Generation   To simulate community language, we instruct Llama-3-8B-Instruct (Touvron et al., 2023) to perform linguistic style transfer while preserving the original content and context. The model takes post titles and comment content as input and generates five variations of each comment representing different scales of normness, such as: “Very Toxic,” “Somewhat Toxic,” “Neutral,” “Somewhat Supportive,” “Very Supportive” for the Toxic–Supportive dimension. See Appendix I.1 for exact prompts used for each norm dimension.
@@
Data Processing   We sample 50K comments per subreddit [^6] The data is sampled from the subset not used to train the community preference predictor, which ensures that the trained CPP model does not perform any inference on its training data in the community preference distillation stage. to use as the seed comments for community language simulation. To ensure the synthetic data quality, we apply preprocessing, lexical, fluency, and content preservation filters (Figure 2) inspired by prior works in style transfer evaluation Briakou et al. (2021); Mir et al. (2019), removing 33% of the synthetic comments. Filter details are reported in Appendix I.2.
Evaluation   Three expert annotators familiar with each topical group evaluated 5 original–synthetic comment pairs per subreddit, resulting in 195 annotated samples. The annotators assessed (1) content similarity of the pair, (2) fluency, (3) authorship (LLM or human), and (4) overall quality (i.e., whether the comment is suitable to be posted in the subreddit) of each comment. Table 2 shows that synthetic data fluently preserves content, and is of good overall quality. Expert annotators failed to identify synthetic data as machine-generated 50% of the time. Moreover, postmortem interviews revealed that being “politically correct” is a strong identifier for machine-ness, and authorship is indistinguishable in science and finance topics. Overall, these results validate the quality of the filtered data. Further details are in Appendix E.2.
@@ 4.3 Community Preference Predictor (CPP)
Data Preprocessing   We take all first-level comments and their associated up-/down-vote counts. We filter out samples that are deleted, edited, created after 1 day of the post creation time, or created within 1 day of the data scraping time as they would skew the true preference.
Models   CPP is fine-tuned on the DialogRPT model—a dialog response ranking GPT-2 based model trained on 133M data from Reddit (Gao et al., 2020). Initializing CPP with DialogRPT weights enhances its understanding of general dialogue dynamics and community preferences. We train a distinct CPP model for each selected subreddit; the fine-tuning process customizes the model to better predict the preference habits of the specific community. See Appendix H.1 for training details.
@@
Evaluation   Following Gao et al. (2020), model performance is evaluated using binary accuracy: whether the relative relations between the predictions and ground truth labels of comment pairs align. We found that including contextual information such as the post title and time of the post significantly improves the accuracy, while adding author information only helps in certain subreddits such as r/libertarian. The most performant setup, comment+post+timestamp, achieved an accuracy of 73.9% (±4.1), suggesting reliable prediction performance. See detailed results in Appendix H.2.
@@ 5 Results
Our RPM results demonstrate how a community’s preferences varies with the scale of normness. We highlight two key RPM plots—formality preferences in politics subreddits and supportiveness preferences in science subreddits—to validate ValueScope in Figure 3 (with full results in Appendix N).
@@
Feedback intensity distinctly varies across topics. Gender-related subreddits extensively reward behaviors aligned with their values, suggesting a strong preference for promoting norms that enhance inclusivity and respect. In contrast, politics subreddits are more conservative with rewards, possibly due to explicit rules against “disproportionate upvoting,” “brigading,” and “upvote spam,” which aim to prevent bias. These regulations may contribute to more measured rewards. Lastly, PRD variations across norm dimensions reveal which normative behaviors are most regulated. The serious–humorous, genuine–sarcastic and concise–verbose dimensions witness the most intense regulation in all topical groups, suggesting the importance of tone and authenticity of interactions in cultivating social identity (Brown, 2022).
@@ 6 Analysis
Our results reveal significant variability in user normative behaviors between the selected subreddit pairs. For example, users in r/wallstreetbets, known for its usage of profane jargon and aggressive trading strategies Herrman (2021), significantly modify their behaviors in r/stocks and r/pennystocks, but adapt much less in the spinoff subreddit r/wallstreetbetsnew. Additionally, user behaviors tend to remain consistent in identity-related subreddits (e.g., r/askwomen, r/askmen) or those with competing relationships (r/republican, r/democrats). These findings highlight the context-specific nature of community norm adaptations by users. We also observe that users are more likely to change their formality to fit different subreddit contexts than other dimensions, such as humor, indicating that certain norms are more malleable and adaptable than others.

### v2

@@ 1 Introduction
Social norms—the perceived, informal, and mostly unwritten rules that govern acceptable behaviors within a community—are foundational to understanding the dynamics of social interactions and shaping the community’s identity UNICEF (2021). Social values, in turn, are the deeper ideals and principles that a community aspires to uphold, guiding the creation and enforcement of these norms (McClintock, 1978). Social norms and values emerge organically through the interplay of behaviors Bicchieri et al. (2023) and are difficult to grasp without gaining experience of the community firsthand. This complexity poses challenges for new users to assimilate Lampe et al. (2014) and makes it difficult for automatic community moderation systems (Park et al., 2021).
@@
    We offer new insights into social dynamics, especially how they evolve over time. These findings have important scientific and practical implications for social scientists and community moderators, helping them identify norms that are likely to change and enabling proactive intervention.
@@ 2 Related Works
Norms and Values in Online Communities   Computational studies have examined linguistic norms and semantic changes in online communities Lucy and Bamman (2021); Del Tredici and Fernández (2018); Kershaw et al. (2016); Danescu-Niculescu-Mizil et al. (2013); Hemphill and Otterbacher (2012); Del Tredici and Fernández (2017); Snoswell et al. (2023); Chancellor et al. (2018). However, these often focus narrowly on language use and neologisms, neglecting the broader spectrum of community values influenced by feedback. Prior research has utilized Schwartz’s Theory of Human Values to estimate values of online communities van der Meer et al. (2023); Borenstein et al. (2024). Weld et al. (2024) has employed survey methods to create a taxonomy of online community values. While some research has addressed explicit governance Chandrasekharan et al. (2018); Fiesler et al. (2018); Park et al. (2021) or qualitatively studied implicit norms Kasunic and Kaufman (2018); Shen and Rosé (2022), our approach fills the gap by (1) focusing on a range of implicit norms (e.g., formality and sarcasm) automatically selected through a generalizable norm induction process, and (2) analyzing collective community preference over behaviors along the selected norm dimensions to capture a comprehensive spectrum of community values, which can provide a more fine-grained and objective measurement for alignment Bergman et al. (2024); Findeis et al. (2024).
@@ 3.1 The ValueScope Framework
Theoretical Background   Community members acquire social adeptness by learning unwritten rules, or implicit norms with feedback from others to guide their behaviors (Coutu, 1951; Zhang et al., 2023). The Return Potential Model (Jackson, 1966, RPM) quantifies these norms by mapping the return potential—expected (dis)approval—across different behaviors. Individuals in a community adjust their actions based on the learned mental model of return potential. We propose ValueScope, a computational framework that adapts RPM to analyze the expected community preference to behaviors with varying normness scales (i.e., conforming to a norm dimension to different extents), offering scalable insights into community values.
Problem Definition   Let $\mathcal{C}$ be communities, $\mathcal{A}$ be comments, and $\mathcal{D}$ be norm dimensions (e.g., sarcasm). For an arbitrary community $c\in\mathcal{C}$ and norm dimension $d\in\mathcal{D}$, ValueScope measures the normness scale $\Phi$ via the Normness Scale Predictor, $\Phi_{d}:\mathcal{A}\rightarrow\mathbb{R}$, and the community preference $\Psi$ via the Community Preference Predictor, $\Psi_{c}:\mathcal{A}\rightarrow\mathbb{R}$, of all $N$ comments in $c$: $\mathcal{A}_{c}$. [^2] Empirically, we perform a distillation step to mitigate confounding factors and distill scores as derived in §3.2 and §3.3—we simply take the delta between two comments $(a_{i},a_{i}^{\prime})$ to get $\nabla\Phi_{d}:(\mathcal{A}\times\mathcal{A})\rightarrow\mathbb{R}=\Phi_{d}(a_{i}^{\prime})-\Phi_{d}(a_{i})$ and $\nabla\Psi_{c}:(\mathcal{A}\times\mathcal{A})\rightarrow\mathbb{R}=\Psi_{c}(a_{i}^{\prime})-\Psi_{c}(a_{i})$. For an arbitrary range of normness scales $\Phi_{d}^{i}:=[\phi_{d}^{\prime},\phi_{d}^{\prime\prime})$ (e.g., “somewhat sarcastic”), we take the set of comments $\mathcal{A}_{c,d}^{i}:=\{a_{i}|\Phi_{d}(a_{i})\in\Phi_{d}^{i}\}$ with normness scales in the given range, and let $N_{c,d}^{i}:=||\mathcal{A}_{c,d}^{i}||$ be the number of comments in this subset. We compute the community preference of these comments:
@@
Interpreting ValueScope   Via ValueScope, we quantitatively observe a number of features of the RPM model proposed in social science literature Jackson (1966); Nolan (2015); Linnan et al. (2005). Specifically, we use the point of maximum return—the highest point on the RPM curve—to locate the ideal normative behavior one should follow to maximize community preference, and the potential return difference—total positive feedback minus total negative feedback—to discover norm regulation strategies; i.e., whether the community tends to use reward or punishment to guide the formation and adaptation of its values.
@@ 3.2 Normness Scale Predictor (NSP)
Normness Distillation   The normness distillation stage addresses two key challenges. First, unlike survey-based social science studies, our approach observes normative behaviors post-hoc, lacking the opportunity to explore “alternative behaviors.” We attempt to recreate the “hypothetical conditions” proposed in Jackson (1966), in which the individual considers alternative options to maximize return (Zhang et al., 2023). We achieve this with a Community Language Simulation (CLS) module, which generates comments identical to the original, except for controlled variations in one norm dimension. This design ensures that any confounding factors are controlled, as the generated comment remains identical to the original except for the intended variation. We then apply the normness measurement module to quantify the normness scales of the transformed comments. E.g., for an original comment, “ty!,” we generate “thank you” by varying formality, and obtain formality scales of $0.2$ and $0.4$, respectively.
@@ 4.2.1 Normness Measurement
Data Preprocessing   Each topical group and norm dimension except for the verbosity dimension  [^4] Instead of training a verbosity scale classifier, we measure verbosity using character count and compute winrates in the range [0-1] based on the count to align with other dimensions. has a dedicated classifier model, enabling comparisons across similar subreddits. Normness measurement relies on synthetic labels generated through stratified sampling and automatic labeling. During the sampling stage, comments are rated on a 5-point Likert scale by GPT-3.5 (Brown et al., 2020) to gauge normness (see Appendix D for the Likert scale details; Appendix E.2 for GPT-3.5 rating evaluation details). Then, 10 comments are sampled per scale point per subreddit, resulting in 150 comments per topic (200 for finance with 4 subreddits included). From these, 1,250 comment pairs are randomly selected to create binary synthetic labels using GPT-4  [^5] We used GPT-3.5 for stratified sampling to save costs, as perfect precision was unnecessary. GPT-4, which performed best in our evaluation (Table 7), assigned high-quality labels to pairwise comments. See Appendix F for GPT cost estimations. (OpenAI et al., 2024); we detail the GPT-4 prompt tuning and synthetic label evaluations in Appendix E.3. We train DeBERTa-base He et al. (2020) with the synthetic labels for each of the 4 topic groups and 5 norm dimensions with training details in Appendix G.1.
Evaluation   To evaluate the quality of GPT-4 generated training labels and the NSP models, we curate a high-quality human annotation set of 450 samples for each norm dimension, where each sample is annotated by 3 annotators with an average inter-annotator agreement, measured by Fleiss’s kappa, of 0.56 (see Appendix E.1 for annotation details). We then compare the GPT-4 generated labels against the human annotations and present the evaluation results in Appendix E.4, with the evaluation of the NSP models detailed in Appendix G.3. Overall, we found that GPT-4 achieved average F1 scores ranging from 75.2-82.4 across the topical groups. In comparison, the NSP models obtained average F1 scores ranging from 74.2-83.0, further validating the quality of the NSP models.
@@ 4.2.2 Community Language Simulation
The norm distillation stage of NSP employs a community language simulation module to synthesize comments and control for norm variations. [^6] We include confounding factor baselines where only original comments with real upvotes are plotted in Appendix N. We found that original comments are unevenly distributed across the normness scales in different subreddits (e.g., r/shittyaskscience is mostly sarcastic, r/askscience is mostly serious), making direct comparison challenging and thus further justifying the need to use the CLS module.
Data Generation   To simulate community language, we instruct Llama-3-8B-Instruct (Touvron et al., 2023) to perform linguistic style transfer while preserving the original content and context. The model takes post titles and comment content as input and generates five variations of each comment representing different normness scales, such as: “Very Toxic,” “Somewhat Toxic,” “Neutral,” “Somewhat Supportive,” “Very Supportive” for the Toxic–Supportive dimension. See Appendix H.1 for the prompts used for each norm dimension.
@@
Data Processing   We sample 50K comments per subreddit [^7] The data is sampled from the subset not used to train the community preference predictor, which ensures that the trained CPP model does not perform any inference on its training data in the community preference distillation stage. to use as the seed comments for community language simulation. To ensure the synthetic data quality, we apply preprocessing, lexical, fluency, and content preservation filters (Figure 2) inspired by prior works in style transfer evaluation Briakou et al. (2021); Mir et al. (2019), removing 33% of the synthetic comments (Appendix H.2).
Evaluation   Three expert annotators familiar with each topical group evaluated 5 original–synthetic comment pairs per subreddit, resulting in 195 annotated samples. The annotators assessed (1) content similarity of the pair, (2) fluency, (3) authorship (LLM or human), and (4) overall quality (i.e., whether the comment is suitable to be posted in the subreddit) of each comment. Table 1 shows that synthetic data fluently preserves content, and is of good overall quality. Expert annotators failed to identify synthetic data as machine-generated 50% of the time. Moreover, postmortem interviews revealed that being “politically correct” is a strong identifier for machine-ness, and authorship is indistinguishable in science and finance topics. Overall, these results validate the quality of the filtered data. Further details are in Appendix H.3.
@@ 4.3 Community Preference Predictor (CPP)
Data Preprocessing   We take all first-level comments and their associated up-/down-vote counts. We exclude comments deleted, edited, created after 1 day of the post creation time, or created within 1 day of data scraping to obtainthe true preference.
Models   CPP is fine-tuned on the DialogRPT model—a dialog response ranking GPT-2 based model trained on 133M data from Reddit (Gao et al., 2020). Initializing CPP with DialogRPT weights enhances its understanding of general dialogue dynamics and community preferences. We train a distinct CPP model for each selected subreddit; the fine-tuning process customizes the model to better predict the preference habits of the specific community. See Appendix I.1 for training details.
@@
Evaluation   Following Gao et al. (2020), model performance is evaluated using binary accuracy: whether the relative relations between the predictions and ground truth labels of comment pairs align. We found that including contextual information such as the post title and time of the post significantly improves the accuracy, while adding author information only helps in certain subreddits such as r/libertarian. The most performant setup, comment+post+timestamp, achieved an accuracy of 73.9% (±4.1), suggesting reliable prediction performance. See detailed results in Appendix I.2.
@@ 5 Results
Our RPM results demonstrate how a community’s preferences varies with the scale of normness. We highlight two key RPM plots—formality preferences in politics subreddits and supportiveness preferences in science subreddits—to validate ValueScope in Figure 3 (with full results in Appendix M).
@@
Feedback intensity distinctly varies across topics. Gender-related subreddits extensively reward behaviors aligned with their values, suggesting a strong preference for promoting norms that enhance inclusivity and respect. Politics subreddits are more conservative with rewards, possibly due to explicit rules against “disproportionate upvoting” and “brigading,” which aim to prevent bias. These regulations may contribute to more measured rewards. Lastly, PRD variations across norm dimensions reveal which normative behaviors are most regulated. The serious–humorous, genuine–sarcastic and concise–verbose dimensions witness the most intense regulation in all groups, suggesting the importance of tone and authenticity of interactions in cultivating social identity (Brown, 2022).
@@ 6 Analysis
Our results reveal significant variability in user normative behaviors between the selected subreddit pairs. For example, users in r/wallstreetbets, known for its usage of profane jargon and aggressive trading strategies Herrman (2021), significantly modify their behaviors in r/stocks and r/pennystocks, but adapt much less in the spinoff subreddit r/wallstreetbetsnew. Additionally, user behaviors tend to remain consistent in identity-related subreddits (e.g., r/askwomen, r/askmen) or those with competing relationships (r/republican, r/democrats), highlighting the context-specific nature of community norm adaptations by users. We also observe that users are more likely to change their formality to fit different subreddit contexts than other dimensions, such as humor, indicating that certain norms are more malleable and adaptable than others.
@@ 7 Conclusion & Future Directions
Acknowledgement
This work was supported by the National Science Foundation (Grant No. IIS-2143529) and the National Institutes of Health (Grant Nos. 5R21DA056725-02, R21DA056725-01A1). We gratefully acknowledge support from the National Science Foundation under CAREER Grant No. IIS2142739, and NSF grants No. IIS2125201 and IIS2203097.